Hydrogen sulfide, H2S, is a flammable gas with an odor like rotten eggs. How many moles of hydrogen sulfide are in 1.47 g of hydrogen sulfide?
1
Expert's answer
2019-09-17T04:59:04-0400

The molar mass of H2S is:


"M(H_2S)=2*M(H)+M(S)=2*1g\/mol+32g\/mol=34g\/mol"


Thus, the amount in moles of 1.47 g of H2S:


"n(H_2S)=\frac{m(H_2S)}{M(H_2S)}=\frac{1.47g}{34g/mol}=0.043mol"